### v1

# The Confidence Paradox: Can LLM Know When It’s Wrong?

Abstract
Document Visual Question Answering (DocVQA) systems are increasingly deployed in real-world applications, yet they remain ethically opaque-often producing overconfident answers to ambiguous questions or failing to communicate uncertainty in a trustworthy manner. This misalignment between model confidence and actual knowledge poses significant risks, particularly in domains requiring ethical accountability. Existing approaches such as LayoutLMv3, UDOP, and DONUT have advanced state-of-the-art performance by focusing on architectural sophistication and accuracy; however, they fall short in ethical responsiveness.
To address these limitations, we introduce HonestVQA, a self-supervised honesty calibration framework for ethically aligned DocVQA. Our model-agnostic method quantifies uncertainty to identify knowledge gaps, aligns model confidence with actual correctness using weighted loss functions, and enforces ethical response behavior via contrastive learning. We further introduce two principled evaluation metrics-Honesty Score (H-Score) and Ethical Confidence Index (ECI)-to benchmark alignment between confidence, accuracy, and ethical communication. Empirically, HonestVQA improves DocVQA accuracy by up to 4.3% and F1 by 4.3% across SpDocVQA, InfographicsVQA, and SROIE datasets. It reduces overconfidence, lowering H-Score and ECI by 0.072 and 0.078, respectively. In cross-domain evaluation, it achieves up to 78.9% accuracy and 76.1% F1-score, demonstrating strong generalization. Ablation shows a 3.8% drop in accuracy without alignment or contrastive loss.
The Confidence Paradox: Can LLM Know When It’s Wrong?
Sahil Tripathi1, Md Tabrez Nafis2, Imran Hussain3, Jiechao Gao4thanks: Corresponding Author: jiechao@stanford.edu 1, 2, 3Jamia Hamdard, New Delhi, India 4Center for SDGC, Stanford University, California, USA
1 Introduction
Document Visual Question Answering (DocVQA) has emerged as a key challenge in multimodal AI Wang et al. (2025), enabling systems to answer questions based on visual and textual content in documents such as invoices, forms, contracts, and academic papers. These systems are widely deployed in enterprise automation Jiang et al. (2024), legal analysis Liu et al. (2024), and assistive technologies Zeng et al. (2024). However, despite their growing utility, DocVQA systems often lack ethical transparency—frequently returning confidently incorrect answers to ambiguous, adversarial, or under-specified queries. For instance, a system may assert the total invoice amount with high confidence even when the relevant table is partially occluded, or confidently misinterpret a scanned signature line as a date. Such failures can propagate serious downstream consequences, including legal misinterpretation, misinformation, or financial misjudgment.
However, the crux of the problem lies in the inability of existing DocVQA models to communicate uncertainty in a calibrated, ethically responsible manner. While State-of-the-Art (SOTA) systems such as LayoutLMv3111https://huggingface.co/microsoft/layoutlmv3-base Fujitake (2024), UDOP [^2] https://huggingface.co/microsoft/udop-large Wang et al. (2023a), and DONUT [^3] https://huggingface.co/naver-clova-ix/donut-base Li et al. (2024) focus on improving accuracy through sophisticated architecture and pretraining strategies, they fall short in aligning model confidence with actual knowledge. LayoutLMv3Fujitake (2024) tends to prioritize exact answers over conveying doubt, UDOP Wang et al. (2023a) frequently errs on the side of over-caution without actionable explanations, and DONUT Li et al. (2024) offers no uncertainty estimation at all—leading to ethically untrustworthy behavior in ambiguous scenarios. Therefore, recent advances in AI alignment research have emphasized the importance of ethical calibration Rao et al. (2023), including honesty Yang et al. (2024), confidence-awareness Stangel et al. (2025), and transparent failure modes Stewart et al. (2023). However, these insights have yet to be meaningfully integrated into DocVQA systems. To address these critical gaps, we propose HonestVQA, a self-supervised framework that calibrates model confidence to reflect its underlying knowledge and ethical responsibility. Our approach is model-agnostic and integrates three key components: (1) uncertainty quantification to identify knowledge gaps, (2) confidence-accuracy alignment through weighted loss optimization, and (3) contrastive learning to enforce ethical response boundaries in ambiguous contexts. We also introduce two novel evaluation metrics: i) Honesty Score (H-Score), which captures the alignment between confidence and correctness, and ii) Ethical Confidence Index (ECI), which evaluates whether high-confidence answers are ethically warranted.
2 Related Work
Recent research has increasingly focused on improving the reliability and interpretability of AI systems, especially in high-stakes domains. While core DocVQA models like have been discussed in Section 1, here we focus on complementary areas that our framework draws from—confidence calibration, ethical modeling, and contrastive learning. Confidence calibration techniques such as temperature scaling Xie et al. (2024) and label smoothing Müller et al. (2019) aim to align predicted probabilities with empirical accuracies. However, these methods are typically post-hoc and task-agnostic, often failing to generalize in multimodal settings. Selective prediction frameworks such as Chen et al. (2023) allow models to abstain from uncertain answers, but they usually rely on fixed thresholds and lack principled mechanisms to model epistemic uncertainty in visually grounded tasks like DocVQA. However, in the area of ethical and honest AI, efforts such as instruction tuning for alignment Zhang et al. (2023) and calibrated language modeling Zhu et al. (2023) emphasize epistemic humility—training models to express uncertainty when appropriate. However, these approaches are primarily developed for language-only models and remain underexplored in multimodal tasks involving structured visual data. Whereas, contrastive learning has shown strong performance in aligning multimodal representations, with frameworks like CLIP Gao et al. (2024) and ALIGN Wang et al. (2023b) leveraging contrastive objectives for image-text alignment. While effective at learning generalizable embeddings, such methods are not designed to enforce ethical boundaries or distinguish between honest and overconfident outputs in ambiguous scenarios.
Algorithm 1 HonestVQA Training Algorithm
1:Pretrained model $f_{\theta}$, input $(D,Q,y^{*})$, thresholds $\delta$, $\tau_{1}$, $\tau_{2}$, weights $\alpha$, $\beta$, $m$, $\lambda_{1}$, $\lambda_{2}$
2:Calibrated DocVQA wrapper
3:Compute $P(y\mid D,Q)\leftarrow f_{\theta}(D,Q)$
4:Compute confidence $\mathcal{C}=\max_{i}P(y_{i})$ and entropy $\mathcal{U}=-\sum_{i}P(y_{i})\log P(y_{i})$
5:Predict $\hat{y}\leftarrow\arg\max_{y}P(y)$
6:$\mathcal{L}_{\text{align}}\leftarrow\alpha\cdot\mathbb{1}[\hat{y}\neq y^{*}]\cdot\mathcal{C}+\beta\cdot\text{CE}(\hat{y},y^{*})$
7:if $\text{WMD}(\hat{y},y^{*})<\delta$ then
8:     $h_{\text{pos}}\leftarrow\text{Embed}(\hat{y})$
9:end if
10:if $\hat{y}\neq y^{*}\land\mathcal{C}>\tau_{1}\land\mathcal{U}<\tau_{2}$ then
11:     $h_{\text{neg}}\leftarrow\text{Embed}(\hat{y})$
12:end if
13:Compute $\mathcal{L}_{\text{contrast}}=\max(0,m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})+\text{sim}(h_{\text{anchor}},h_{\text{neg}}))$
14:$\mathcal{L}_{\text{total}}\leftarrow\lambda_{1}\cdot\mathcal{L}_{\text{align}}+\lambda_{2}\cdot\mathcal{L}_{\text{contrast}}$
15:Update projection head using $\mathcal{L}_{\text{total}}$
3 Methodology
As discussed earlier, HonestVQA is a model-agnostic calibration framework designed to enhance ethical transparency in DocVQA systems. It operates as a wrapper around pretrained DocVQA models (in our work, we evaluate our framework on top of pretrained models such as LayoutLMv3Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024) to demonstrate its generalizability), injecting uncertainty-aware alignment and contrastive reasoning to reduce overconfident yet incorrect outputs. The broader process of the HonestVQA is illustrated in Algorithm 1.
3.1 Uncertainty Quantification Module
Given a document $D$ and a question $Q$, we use a pretrained DocVQA model $f_{\theta}$ that maps $(D,Q)$ to an answer distribution $P(y\mid D,Q;\theta)$. To quantify the model’s epistemic uncertainty, we compute the softmax entropy of the output distribution according to Equation (1).
$\mathcal{U}(D,Q)=-\sum_{i=1}^{|Y|}P(y_{i}\mid D,Q)\log P(y_{i}\mid D,Q)$
Here, $|Y|$ denotes the size of the answer space. Higher entropy corresponds to greater uncertainty. We also define a maximum-confidence score as shown in Equation (2).
$\mathcal{C}(D,Q)=\max_{i}P(y_{i}\mid D,Q)$
This dual view captures both dispersion and peakiness in the output distribution. Following recent work Pearce et al. (2021), we identify overconfident failure cases as those where $\mathcal{C}(D,Q)$ is high despite $\mathcal{U}(D,Q)$ being non-negligible. These metrics are computed during training and inference, and $\mathcal{U}(D,Q)$ serves as a routing signal for sampling in the contrastive module, though it is not explicitly penalized.
3.2 Confidence-Accuracy Alignment Module
To align the model’s confidence with its accuracy, we introduce a calibration-aware loss that penalizes incorrect predictions more strongly when made with high confidence. Let $\hat{y}$ denote the predicted answer and $y^{*}$ the ground truth. We define the alignment loss according to Equation (2).
$\mathcal{L}_{\text{align}}=\alpha\cdot\mathbb{1}[\hat{y}\neq y^{*}]\cdot\mathcal{C}(D,Q)+\beta\cdot\text{CE}(\hat{y},y^{*})$
Here, CE is the standard cross-entropy loss. Hyperparameters $\alpha$ and $\beta$ control the influence of confidence-penalization and prediction error, respectively.
3.3 Contrastive Ethical Enforcement Module
To further refine the model’s response space under ambiguity, we introduce a contrastive loss that structurally separates ethically misaligned or misleading answers from calibrated, semantically valid responses. Given a query-answer embedding $h_{\text{anchor}}$, we identify a positive sample $h_{\text{pos}}$ (semantically similar and ethically aligned) and a negative sample $h_{\text{neg}}$ (incorrect, overconfident, or potentially misleading). The contrastive loss is defined as according to Equation (3).
$\mathcal{L}_{\text{contrast}}=$ $\max\Big{(}0,\ m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})$
$+\ \text{sim}(h_{\text{anchor}},h_{\text{neg}})\Big{)}$ (3)
where $\text{sim}(\cdot)$ denotes cosine similarity, and $m$ is a margin hyperparameter. We use a projection head atop the DocVQA encoder to map answer embeddings into a low-dimensional calibrated honesty space. Where, positive pairs are identified using a combination of Word Mover’s Distance (WMD) and agreement with ground truth as shown in Equations (4), and (5), where $\delta$ is a tunable similarity threshold.
$\text{WMD}(\hat{y},y^{*})<\delta$
$\hat{y}\in\mathcal{A}_{\text{aligned}}\Longrightarrow$ semantically valid
and agrees with ground truth. (5)
Whereas, negative samples are drawn from high-confidence as shown in Equation (6).
$\hat{y}_{\text{neg}}:$ $\mathbb{1}[\hat{y}_{\text{neg}}\neq y^{*}]$
$\land\ \mathcal{C}(D,Q)>\tau_{1}$
$\land\ \mathcal{U}(D,Q)<\tau_{2}$ (6)
where $\tau_{1}$ and $\tau_{2}$ are confidence and entropy thresholds, respectively.
3.4 Training Module
The overall training loss combines the alignment and contrastive objectives according to Equation (7).
$\mathcal{L}_{\text{total}}=\lambda_{1}\cdot\mathcal{L}_{\text{align}}+\lambda_{2}\cdot\mathcal{L}_{\text{contrast}}$
Here, $\lambda_{1}$ and $\lambda_{2}$ control the relative weight of each loss term. Training is conducted end-to-end using batches sampled from standard DocVQA datasets, where each sample includes contrastive triplets and confidence-aware labels. The projection head is fine-tuned during training, while the base DocVQA encoder remains frozen. At inference, $\mathcal{C}(D,Q)$ and $\mathcal{U}(D,Q)$ may be used to suppress or abstain from answering under high uncertainty. Note: In this work, we define ethical calibration as the act of reducing confidently incorrect answers, especially under ambiguity or lack of sufficient visual-textual grounding. HonestVQA is not a moral arbiter but a mechanism for promoting caution and transparency in DocVQA behavior.
4 Experimental Setup
4.1 Datasets
We evaluate HonestVQA on three diverse and challenging datasets: SpDocVQA Mathew et al. (2020), InfographicsVQA Mathew et al. (2022), and SROIE [^4] https://rrc.cvc.uab.es/?ch=17&com=downloads. SpDocVQA comprises multilingual scanned documents requiring structured comprehension and spatial reasoning. InfographicsVQA presents visually dense infographic images with complex layouts and multi-modal reasoning demands. SROIE is an entity-level extraction dataset involving semi-structured receipts, demanding high accuracy and ethical response handling due to potential financial implications. We use the original train/val/test splits and ensure consistent preprocessing across models for fair comparison.
4.2 Evaluation Metrics
To comprehensively assess the performance and ethical alignment of our framework, we employ standard accuracy metrics alongside two novel measures designed to evaluate calibration and honesty. First, we report conventional Accuracy and F1 scores (i.e. macro) to quantify answer correctness. To capture the alignment between model confidence and actual correctness, we introduce the Honesty Score (H-Score), which penalizes overconfident incorrect predictions while rewarding calibrated confidence on correct answers. Additionally, the Ethical Confidence Index (ECI) evaluates the model’s ability to appropriately express uncertainty, especially under ambiguous or insufficient information. Note: In the tables, color coding is used to indicate the performance of the models: Blue represents the best performance, Green indicates the second-best performance, and Yellow denotes the third-best performance. Whereas, ↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
4.2.1 Theoretical Guarantees for Evaluation Metrics
In this section, we provide formal lemmas and proofs to establish the theoretical soundness of the proposed Honesty Score (H-Score) and Ethical Confidence Index (ECI), which measure the alignment between model confidence, accuracy, and ethical transparency in DocVQA.
Lemma 4.1 (Calibration Bound of Honesty Score).
Let $C(D,Q)$ be the confidence score output by a DocVQA model for a given document-question pair $(D,Q)$, and let $A(D,Q)\in\{0,1\}$ be the corresponding accuracy indicator, where $1$ denotes a correct answer and 00 an incorrect one. Assume that $C(D,Q)$ is bounded in $[0,1]$. Then, define the Honesty Score $\mathrm{H}$ as according to Equation (8).
$\mathrm{H}=1-\mathbb{E}_{(D,Q)\sim\mathcal{D}}\big{[}|C(D,Q)-A(D,Q)|\big{]}$
where $\mathcal{D}$ is the data distribution. Whereas, $\mathrm{H}$ upper-bounds the expected calibration error between confidence and accuracy according to Equation (9).
$\mathbb{E}_{(D,Q)\sim\mathcal{D}}\big{[}|C(D,Q)-A(D,Q)|\big{]}=1-\mathrm{H}$
Thus, a higher $\mathrm{H}$ implies tighter calibration, indicating fewer overconfident incorrect predictions.
Proof.
By definition, calibration error measures the absolute difference between predicted confidence and true correctness. Since $A(D,Q)$ is binary, the expectation of $|C-A|$ captures the average misalignment. Rearranging, $\mathrm{H}=1-\mathbb{E}[|C-A|]$. Because $|C-A|\in[0,1]$, $\mathrm{H}\in[0,1]$ and is maximized when confidence perfectly matches accuracy. Hence, $\mathrm{H}$ is a valid measure of calibration that upper-bounds expected miscalibration. ∎
Lemma 4.2 (Discriminative Power of Ethical Confidence Index).
Let $\mathcal{C}_{\text{correct}}$ and $\mathcal{C}_{\text{incorrect}}$ denote the random variables corresponding to confidence scores on correctly and incorrectly answered samples respectively. Then, define the Ethical Confidence Index (ECI) as according to Equation (10) which measures the probability that the model assigns higher confidence to correct answers than to incorrect answers.
$\mathrm{ECI}=\mathbb{P}\big{(}\mathcal{C}_{\text{correct}}>\mathcal{C}_{\text{incorrect}}\big{)}$
If the distributions of $\mathcal{C}_{\text{correct}}$ and $\mathcal{C}_{\text{incorrect}}$ are well-separated, i.e., there exists $\epsilon>0$ such that it is defined as according to Equation (11), then $\mathrm{ECI}\geq 1-\epsilon$ indicating strong ethical confidence discrimination.
$\mathbb{P}(\mathcal{C}_{\text{correct}}\leq\mathcal{C}_{\text{incorrect}})<\epsilon$
Proof.
The ECI corresponds exactly to the Area Under the ROC Curve (AUC) when viewing confidence as a score discriminating correct from incorrect answers. By definition, $\mathrm{ECI}=\mathbb{P}(\mathcal{C}_{\text{correct}}>\mathcal{C}_{\text{incorrect}})$. If the two confidence score distributions have minimal overlap (i.e., are well-separated), the probability of $\mathcal{C}_{\text{correct}}\leq\mathcal{C}_{\text{incorrect}}$ is bounded above by a small $\epsilon$. Hence, $\mathrm{ECI}=1-\mathbb{P}(\mathcal{C}_{\text{correct}}\leq\mathcal{C}_{\text{incorrect}})\geq 1-\epsilon$. Thus, a high ECI value indicates that the model reliably assigns higher confidence to correct answers, promoting ethical transparency. ∎
Model SpDocVQA InfographicsVQA SROIE
Accuracy (%) ↑ Macro F1 (%) ↑ Accuracy (%) ↑ Macro F1 (%) ↑ Accuracy (%) ↑ Macro F1 (%) ↑
Base Models
LayoutLMv3 (base) Fujitake (2024) 72.3 68.5 65.4 62.1 70.0 66.8
UDOP (base) Wang et al. (2023a) 69.7 66.1 62.8 60.0 68.2 64.0
DONUT (base) Li et al. (2024) 70.1 67.0 63.5 60.9 69.0 65.2
With HonestVQA
LayoutLMv3 Fujitake (2024) + HonestVQA 75.9 72.8 69.7 66.3 73.4 70.1
UDOP Wang et al. (2023a) + HonestVQA 73.2 69.4 67.3 63.8 71.0 67.5
DONUT Li et al. (2024) + HonestVQA 74.0 70.5 68.0 64.7 72.2 68.8
Table 1: Answer correctness comparison on SpDocVQA, InfographicsVQA, and SROIE datasets
Model SpDocVQA InfographicsVQA SROIE
H-Score ↓ ECI ↓ H-Score ↓ ECI ↓ H-Score ↓ ECI ↓
Base Models
LayoutLMv3 (base) Fujitake (2024) 0.185 0.210 0.192 0.215 0.188 0.213
UDOP (base) Wang et al. (2023a) 0.198 0.224 0.203 0.230 0.200 0.228
DONUT (base) Li et al. (2024) 0.190 0.218 0.195 0.222 0.192 0.220
With HonestVQA
LayoutLMv3 Fujitake (2024) + HonestVQA 0.113 0.132 0.118 0.138 0.115 0.136
UDOP Wang et al. (2023a) + HonestVQA 0.127 0.147 0.132 0.153 0.129 0.150
DONUT Li et al. (2024) + HonestVQA 0.120 0.139 0.125 0.143 0.122 0.141
Table 2: Calibration metrics on SpDocVQA, InfographicsVQA, and SROIE datasets
4.3 Hyperparameters
The HonestVQA framework employs several key hyperparameters to balance confidence calibration and contrastive learning effectively. We set the confidence penalty weight $\alpha$ to 1.0 and the cross-entropy weight $\beta$ to 0.5 to emphasize penalizing overconfident incorrect predictions while maintaining prediction accuracy. The contrastive margin $m$ is fixed at 0.3 to enforce a moderate separation between positive and negative embeddings. The alignment and contrastive losses are weighted by $\lambda_{1}=1.0$ and $\lambda_{2}=0.7$, respectively, reflecting a slightly stronger emphasis on alignment. For sample selection in the contrastive module, the WMD threshold $\delta$ is set to 0.4, while the confidence and entropy thresholds, $\tau_{1}$ and $\tau_{2}$, are chosen as 0.8 and 0.5, respectively, to effectively identify semantically valid positive samples and high-confidence misleading negatives. Note: We will make the code publicly available post-acceptance.
5 Experimental Analysis
5.1 Comparison with Baselines
We evaluate the effectiveness of HonestVQA by measuring both standard answer correctness metrics and calibration-specific metrics to provide a comprehensive assessment of the model’s performance. Specifically, we compare the base DocVQA models—LayoutLMv3 Fujitake (2024), UDOP Wang et al. (2023a), and DONUT Li et al. (2024)—with their corresponding versions enhanced by the HonestVQA calibration framework. Table 1 presents these results on the three datasets. It is evident that HonestVQA consistently improves accuracy by approximately 3 to 4 percentage points and macro F1-score by nearly 4 points across all base models. For instance, LayoutLMv3 Fujitake (2024) improves from 72.3% to 75.9% in accuracy and from 68.5% to 72.8% in macro F1-score. Similar trends hold for UDOP Wang et al. (2023a) and DONUT Li et al. (2024) models, underscoring the robustness of our approach in enhancing answer correctness. Whereas, Table 2 displays these calibration-specific results for the same set of models and dataset. Notably, the base models exhibit relatively high H-Score and ECI values, indicating frequent instances of unjustified overconfidence. Incorporation of HonestVQA substantially lowers these values, with H-Score decreasing by over 35% and ECI by nearly 40% on average. For instance, LayoutLMv3’s Fujitake (2024) H-Score drops from 0.185 to 0.113, and ECI decreases from 0.210 to 0.132 after calibration. This demonstrates that HonestVQA effectively mitigates the risk of misleading the user by suppressing confident but incorrect answers.
Train Domain Test Domain Model Accuracy (%) ↑ Macro F1 (%) ↑
SpDocVQA InfographicsVQA HonestVQA 74.2 71.8
InfographicsVQA SpDocVQA HonestVQA 78.9 76.1
SROIE InfographicsVQA HonestVQA 70.5 67.2
SpDocVQA SROIE HonestVQA 72.6 69.8
InfographicsVQA SROIE HonestVQA 73.1 70.4
SROIE SpDocVQA HonestVQA 75.0 72.3
Table 3: Cross-domain hallucination detection performance of HonestVQA
Configuration SpDocVQA InfographicsVQA SROIE
Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓
Full HonestVQA Model 75.9 0.113 68.3 0.134 80.2 0.096
No Alignment Loss 72.1 0.172 65.0 0.193 76.4 0.141
No Contrastive Loss 73.0 0.160 65.2 0.189 77.2 0.133
Table 4: Ablation results across three datasets showing the impact of disabling individual components of HonestVQA
5.2 Cross-Domain Generalization Testing
To assess the robustness and generalization ability of HonestVQA framework, we conduct a series of cross-domain testing experiments. These experiments evaluate whether the hallucination detection model trained on one dataset can effectively identify hallucinations in DocVQA outputs on different datasets. Such generalization is crucial in real-world scenarios where the distribution of questions and visual-textual content varies significantly across domains such as scanned documents, infographics, and structurally diverse textual scenes. We train HonestVQA on the source dataset and evaluate it on a different target dataset without any further fine-tuning, measuring hallucination detection performance using Accuracy and F1-score as the primary metrics. From Table 3, we observe that HonestVQA generalizes robustly across domain shifts. Notably, when trained on InfographicsVQA and evaluated on SpDocVQA, the model achieves a high F1-score of 76.1%, outperforming the reverse setting (SpDocVQA $\rightarrow$ InfographicsVQA), which yields 71.8%. This suggests that the high-density, information-rich visual patterns in InfographicsVQA provide transferable inductive biases that enhance hallucination detection in more structured domains like documents. Similarly, the SROIE $\rightarrow$ SpDocVQA setup results in an F1 of 72.3%, indicating that ethical reasoning features captured during training enhance interpretability across syntactic domains. However, the model exhibits relatively lower performance when transferring from SpDocVQA to SROIE (F1 = 69.8%), highlighting the challenges posed by ethical hallucination detection under unfamiliar structural constraints. Nonetheless, the use of uncertainty-aware calibration via confidence-alignment and contrastive ethical loss contributes to soft regularization of decision boundaries, allowing for improved generalization even in low-overlap semantic settings. Interestingly, models trained on SROIE also generalize well to visually and semantically distinct domains such as InfographicsVQA, achieving 67.2% F1. This supports our hypothesis that the inclusion of contrastive ethical supervision enforces more generalizable representations. Furthermore, the relative drop in performance in domain transfer settings (typically within 4–6% of in-domain results) underscores the importance of calibration-aware models in mitigating performance degradation due to domain shift.
Hyperparameter Values Tested SpDocVQA (Acc ↑, H ↓) InfographicsVQA (Acc ↑, H ↓) SROIE (Acc ↑, H ↓)
Alignment Weight $\alpha$ 0.1 (70.5, 0.195) (63.7, 0.211) (74.5, 0.157)
0.5 (74.8, 0.119) (66.9, 0.148) (78.2, 0.112)
1.0 (75.9, 0.113) (68.3, 0.134) (80.2, 0.096)
1.5 (74.6, 0.118) (67.5, 0.141) (79.4, 0.107)
2.0 (72.7, 0.145) (65.8, 0.174) (77.0, 0.134)
Contrastive Margin $m$ 0.1 (71.2, 0.178) (64.5, 0.198) (75.3, 0.151)
0.3 (74.9, 0.120) (66.7, 0.150) (78.7, 0.108)
0.5 (75.9, 0.113) (68.3, 0.134) (80.2, 0.096)
0.7 (74.7, 0.117) (67.1, 0.143) (79.3, 0.105)
1.0 (72.9, 0.153) (65.2, 0.177) (77.2, 0.127)
Loss Weight $\lambda_{1}$ (Alignment) 0.01 (72.3, 0.164) (64.9, 0.192) (75.8, 0.149)
0.05 (74.5, 0.125) (66.5, 0.153) (78.9, 0.109)
0.10 (75.9, 0.113) (68.3, 0.134) (80.2, 0.096)
0.15 (74.2, 0.130) (67.4, 0.142) (79.1, 0.104)
0.20 (72.5, 0.148) (65.9, 0.169) (77.4, 0.129)
Loss Weight $\lambda_{2}$ (Contrastive) 0.01 (72.8, 0.160) (65.0, 0.186) (76.6, 0.141)
0.03 (74.6, 0.122) (66.7, 0.151) (78.8, 0.110)
0.05 (75.9, 0.113) (68.3, 0.134) (80.2, 0.096)
0.07 (74.3, 0.127) (67.0, 0.144) (79.2, 0.103)
0.10 (73.1, 0.142) (65.4, 0.172) (77.3, 0.125)
Table 5: Sensitivity analysis of alignment and contrastive hyperparameters on SpDocVQA, InfographicsVQA, and SROIE. Note: Accuracy (Acc) and H-Score (H) are used as abbreviated. Bold rows indicate optimal settings.
Model SpDocVQA InfographicsVQA SROIE
IoU (%) ↑ Hallucination Acc. (%) ↑ IoU (%) ↑ Hallucination Acc. (%) ↑ IoU (%) ↑ Hallucination Acc. (%) ↑
LayoutLMv3 Fujitake (2024) 58.3 71.2 54.9 69.5 52.1 67.8
DONUT Li et al. (2024) 60.7 73.0 56.8 70.3 53.7 68.4
HonestVQA 69.1 78.5 65.4 76.8 62.7 74.2
Table 6: Multi-modal consistency evaluation across datasets. Note: IoU measures alignment between visual attention and textual grounding. Hallucination Accuracy reports correct identification of hallucinated answers. HonestVQA achieves superior multi-modal alignment and hallucination detection performance.
5.3 Ablation Study
To thoroughly evaluate the contribution of individual modules in HonestVQA, we conduct an ablation study across three datasets by systematically disabling the confidence-accuracy alignment loss and the contrastive ethical enforcement loss. Table 4 shows that the removal of either module consistently degrades performance across all datasets in terms of accuracy and H-Score, confirming their complementary roles. For instance, on SpDocVQA, the full model achieves 75.9% accuracy and an H-Score of 0.113. Removing the alignment loss drops accuracy to 72.1% and worsens H-Score to 0.172. On InfographicsVQA, the full model yields 68.3% accuracy and an H-Score of 0.134, whereas removing contrastive enforcement lowers accuracy to 65.2% and degrades H-Score to 0.189. Similar trends are observed on SROIE, where the full model achieves 80.2% accuracy and 0.096 H-Score, significantly outperforming the ablated variants. We further conduct a hyperparameter sensitivity analysis on alignment weight $\alpha$, contrastive margin $m$, and loss weights $\lambda_{1}$ and $\lambda_{2}$ across three datasets and summarize the trends in Table 5. The model maintains high performance for $\alpha$ between 0.5 and 1.5, margin $m$ from 0.3 to 0.7, and loss weights $\lambda_{1}=0.1$, $\lambda_{2}=0.05$. Deviations outside these ranges result in decreased accuracy or calibration degradation.
Model Inference Time (ms) FLOPs (Giga) Memory Usage (MB)
Avg ↓ Std Dev ↓ Avg ↓ Std Dev ↓ Avg ↓ Std Dev ↓
LayoutLMv3 Fujitake (2024) 112.4 5.1 64.8 1.3 2950 120
UDOP Wang et al. (2023a) 98.7 4.3 58.6 1.1 2710 105
DONUT Li et al. (2024) 105.1 4.7 62.2 1.2 2830 110
HonestVQA 119.6 5.6 69.4 1.4 3075 130
Table 7: Latency and efficiency comparison of HonestVQA and baselines on SpDocVQA, InfographicsVQA, and SROIE datasets. Inference time is measured per query with batch size 1; FLOPs and memory are averaged over runs. HonestVQA incurs a moderate increase in computational cost due to calibration modules but remains practical for deployment.
6 Additional Analysis
6.1 Multimodal Consistency Evaluation
A critical aspect of hallucination detection in DocVQA is the model’s ability to ensure consistency between the visual features and the textual grounding of answers. To evaluate how well HonestVQA aligns the visual and textual modalities in its hallucination judgments, we conduct a multi-modal consistency evaluation across the three datasets. Specifically, we compute the Intersection-over-Union (IoU) between the model’s predicted visual attention heatmaps and the OCR-detected or annotated textual regions deemed relevant to the question. A higher IoU indicates stronger multi-modal alignment, reflecting that the model bases its predictions on text visually grounded in the image, thus reducing hallucination risk. We report the average IoU scores alongside hallucination detection accuracy for HonestVQA and compare it against two strong baselines: LayoutLMv3Fujitake (2024) and DONUT Li et al. (2024) without calibration. The results are summarized in Table 6. As seen in Table 6, HonestVQA consistently achieves significantly higher IoU scores compared to baselines, demonstrating a stronger alignment between the visual evidence and textual regions considered during inference. For instance, on the SpDocVQA dataset, HonestVQA attains an IoU of 69.1%, which is approximately 8.4% absolute improvement over DONUT Li et al. (2024) and 10.8% over LayoutLMv3 Fujitake (2024). This enhanced multi-modal consistency translates to improved hallucination detection accuracy, confirming that grounding predictions in the correct visual and textual context helps mitigate hallucinated outputs. We further analyze the distribution of IoU scores at the instance level and observe that HonestVQA reduces instances with low cross-modal agreement (IoU < 40%) by over 25% relative to the baselines. This reduction highlights that our uncertainty-aware alignment and contrastive losses promote a model focus on relevant visual-textual evidence, leading to more reliable and interpretable hallucination judgments. Note: UDOP Wang et al. (2023a) was excluded from the multi-modal consistency evaluation as it does not provide explicit or interpretable visual attention maps tied to OCR-detected regions, which are essential for computing IoU-based alignment metrics. Unlike LayoutLMv3 Fujitake (2024), DONUT Li et al. (2024), and HonestVQA, which utilize structured visual-textual grounding mechanisms, UDOP Wang et al. (2023a) primarily relies on unified vision-language pretraining without fine-grained token-region correspondence. As a result, evaluating multi-modal alignment using IoU would not be meaningful or comparable for UDOP Wang et al. (2023a).
6.2 Computational Analysis
Practical deployment of DocVQA systems, particularly those incorporating hallucination detection like HonestVQA, requires a thorough understanding of computational costs and resource demands. To assess the efficiency of HonestVQA relative to baseline models, we conduct a detailed latency and resource usage analysis focusing on inference time, Floating-Point Operations (FLOPs), and memory consumption. All models are tested on identical hardware setups: an NVIDIA RTX 3090 GPU and an Intel Xeon CPU, with batch size fixed to 1 to simulate single-query inference typical of interactive systems. Table 7 summarizes the key efficiency metrics averaged over 1000 inference runs per dataset. From Table 7, HonestVQA demonstrates an average inference latency of approximately 119.6 ms per query, which is about 6%-20% slower than the baseline models due to the additional uncertainty calibration and contrastive enforcement modules integrated into the architecture. Despite this, the latency remains within acceptable limits for real-time applications, especially in contexts where reliability and ethical correctness outweigh marginal speed gains. Regarding computational complexity, HonestVQA requires on average 69.4 GFLOPs, representing a 7%-18% increase relative to baselines. This overhead is primarily attributable to the dual loss functions enforcing confidence-accuracy alignment and ethical contrastive constraints, which necessitate additional forward and backward computations during inference calibration. Memory usage also increases modestly to 3075 MB on average, about 5%-14% more than competing models. The memory footprint remains manageable on modern GPUs, supporting the feasibility of deploying HonestVQA on standard deep learning infrastructure without requiring specialized hardware.
7 Conclusion and Future Works
In this work, we introduced HonestVQA, a novel framework that integrates uncertainty-aware alignment and contrastive ethical enforcement to effectively detect hallucinations in DocVQA systems. Through comprehensive experiments on diverse datasets—we demonstrated significant improvements in answer correctness, calibration, and cross-domain generalization compared to strong baselines. Our ablation studies confirmed the complementary role of each component, and efficiency analyses showed HonestVQA practical feasibility. Future research will focus on advancing domain adaptation techniques to further enhance robustness across unseen data distributions, and exploring lightweight calibration modules for deployment on edge devices. Additionally, incorporating user feedback for interactive hallucination correction and extending the framework to multimodal dialogue systems represent promising directions to improve DocVQA reliability and ethical safety.
Limitations
While HonestVQA significantly improves hallucination detection and ethical calibration, several limitations remain. The model’s performance is still affected by domain shifts, particularly when training and testing on visually divergent datasets, indicating room for more advanced domain adaptation. HonestVQA calibration modules introduce additional computational overhead, which may constrain deployment in highly resource-limited environments. Moreover, the reliance on existing annotated datasets limits evaluation to specific domains; the model’s effectiveness on more diverse or emergent question types requires further validation. Finally, although the framework mitigates hallucinations, it does not guarantee complete elimination, highlighting the need for complementary human-in-the-loop verification for critical applications.
Ethics Statement
This work aims to enhance the trustworthiness and ethical reliability of DocVQA systems by reducing hallucinated and potentially misleading answers. HonestVQA promotes transparency through uncertainty-aware calibration, encouraging responsible AI deployment. We acknowledge the risk that no model can be entirely free of errors or biases, especially when applied across diverse real-world scenarios. Thus, we emphasize that HonestVQA is intended as a tool to assist, not replace, human judgment, particularly in high-stakes contexts. All datasets used comply with their respective licenses, and no private or sensitive data was involved. We encourage further research on fairness, bias mitigation, and inclusivity to ensure equitable AI systems, and advocate for ongoing monitoring of model outputs to safeguard against misuse or unintended harm.

### v2

@@
 Abstract
-Document Visual Question Answering (DocVQA) systems are increasingly deployed in real-world applications, yet they remain ethically opaque-often producing overconfident answers to ambiguous questions or failing to communicate uncertainty in a trustworthy manner.
-This misalignment between model confidence and actual knowledge poses significant risks, particularly in domains requiring ethical accountability.
-Existing approaches such as LayoutLMv3, UDOP, and DONUT have advanced state-of-the-art performance by focusing on architectural sophistication and accuracy; however, they fall short in ethical responsiveness.
-To address these limitations, we introduce HonestVQA, a self-supervised honesty calibration framework for ethically aligned DocVQA.
-Our model-agnostic method quantifies uncertainty to identify knowledge gaps, aligns model confidence with actual correctness using weighted loss functions, and enforces ethical response behavior via contrastive learning.
-We further introduce two principled evaluation metrics-Honesty Score (H-Score) and Ethical Confidence Index (ECI)-to benchmark alignment between confidence, accuracy, and ethical communication.
-Empirically, HonestVQA improves DocVQA accuracy by up to 4.3% and F1 by 4.3% across SpDocVQA, InfographicsVQA, and SROIE datasets.
-It reduces overconfidence, lowering H-Score and ECI by 0.072 and 0.078, respectively.
-In cross-domain evaluation, it achieves up to 78.9% accuracy and 76.1% F1-score, demonstrating strong generalization.
-Ablation shows a 3.8% drop in accuracy without alignment or contrastive loss.
+Document Visual Question Answering (DocVQA) models often produce overconfident or ethically misaligned responses, especially under uncertainty.
+Existing models like LayoutLMv3, UDOP, and DONUT focus on accuracy but lack ethical calibration.
+We propose HonestVQA, a model-agnostic, self-supervised framework that aligns model confidence with correctness using weighted loss and contrastive learning.
+We introduce two new metrics—Honesty [^1] In our work, honesty is operationalized as the alignment between model confidence and correctness, with the goal of reducing confidently wrong answers.
+Score (H-Score) and Ethical Confidence Index (ECI)—to evaluate ethical alignment.
+HonestVQA improves accuracy and F1 by up to 4.3% across SpDocVQA, InfographicsVQA, and SROIE datasets, while reducing overconfidence.
+It also generalizes well across domains, achieving 78.9% accuracy and 76.1% F1-score.
 The Confidence Paradox: Can LLM Know When It’s Wrong?
-Sahil Tripathi1, Md Tabrez Nafis2, Imran Hussain3, Jiechao Gao4thanks: Corresponding Author: jiechao@stanford.edu 1, 2, 3Jamia Hamdard, New Delhi, India 4Center for SDGC, Stanford University, California, USA
+Sahil Tripathi1, Md Tabrez Nafis [^1] footnotemark: 1, Imran Hussain1, Jiechao Gao2thanks: Corresponding Author: jiechao@stanford.edu, tabrez.nafis@gmail.com 1Jamia Hamdard, New Delhi, India 2Center for SDGC, Stanford University, California, USA
 1 Introduction
 Document Visual Question Answering (DocVQA) has emerged as a key challenge in multimodal AI Wang et al.
-(2025), enabling systems to answer questions based on visual and textual content in documents such as invoices, forms, contracts, and academic papers.
-These systems are widely deployed in enterprise automation Jiang et al.
+(2025), enabling models to answer questions based on visual and textual content in documents such as invoices, forms, contracts, and academic papers.
+These models are widely deployed in enterprise automation Jiang et al.
 (2024), legal analysis Liu et al.
@@
 (2025).
-However, despite their growing utility, DocVQA systems often lack ethical transparency—frequently returning confidently incorrect answers to ambiguous, adversarial, or under-specified queries.
-For instance, a system may assert the total invoice amount with high confidence even when the relevant table is partially occluded, or confidently misinterpret a scanned signature line as a date.
+However, despite their growing utility, DocVQA models often lack ethical transparency—frequently returning confidently incorrect answers to ambiguous, adversarial, or under-specified queries.
+For instance, a model may assert the total invoice amount with high confidence even when the relevant table is partially occluded, or confidently misinterpret a scanned signature line as a date.
 Such failures can propagate serious downstream consequences, including legal misinterpretation, misinformation, or financial misjudgment.
 However, the crux of the problem lies in the inability of existing DocVQA models to communicate uncertainty in a calibrated, ethically responsible manner.
-While State-of-the-Art (SOTA) systems such as LayoutLMv3111https://huggingface.co/microsoft/layoutlmv3-base Fujitake (2024), UDOP [^2] https://huggingface.co/microsoft/udop-large Wang et al.
+While State-of-the-Art (SOTA) models such as LayoutLMv3222https://huggingface.co/microsoft/layoutlmv3-base Fujitake (2024), UDOP [^3] https://huggingface.co/microsoft/udop-large Wang et al.
 (2023a), and DONUT [^4] https://huggingface.co/naver-clova-ix/donut-base Li et al.
@@
 (2023).
-However, these insights have yet to be meaningfully integrated into DocVQA systems.
+However, these insights have yet to be meaningfully integrated into DocVQA models.
 To address these critical gaps, we propose HonestVQA, a self-supervised framework that calibrates model confidence to reflect its underlying knowledge and ethical responsibility.
@@
 2 Related Work
-Recent research has increasingly focused on improving the reliability and interpretability of AI systems, especially in high-stakes domains.
+Recent research has increasingly focused on improving the reliability and interpretability of AI models, especially in high-stakes domains.
 While core DocVQA models like have been discussed in Section 1, here we focus on complementary areas that our framework draws from—confidence calibration, ethical modeling, and contrastive learning.
-Confidence calibration techniques such as temperature scaling Xie et al.
+Confidence calibration models such as temperature scaling Xie et al.
 (2024) and label smoothing Müller et al.
 (2019) aim to align predicted probabilities with empirical accuracies.
-However, these methods are typically post-hoc and task-agnostic, often failing to generalize in multimodal settings.
+However, these models are typically post-hoc and task-agnostic, often failing to generalize in multimodal settings.
 Selective prediction frameworks such as Chen et al.
@@
 (2023) emphasize epistemic humility—training models to express uncertainty when appropriate.
-However, these approaches are primarily developed for language-only models and remain underexplored in multimodal tasks involving structured visual data.
+However, these researches are primarily developed for language-only models and remain underexplored in multimodal tasks involving structured visual data.
 Whereas, contrastive learning has shown strong performance in aligning multimodal representations, with frameworks like CLIP Gao et al.
@@
 (2023b) leveraging contrastive objectives for image-text alignment.
-While effective at learning generalizable embeddings, such methods are not designed to enforce ethical boundaries or distinguish between honest and overconfident outputs in ambiguous scenarios.
+While effective at learning generalizable embeddings, such models are not designed to enforce ethical boundaries or distinguish between honest and overconfident outputs in ambiguous scenarios.
 Algorithm 1 HonestVQA Training Algorithm
@@
 5:Predict $\hat{y}\leftarrow\arg\max_{y}P(y)$
-6:$\mathcal{L}_{\text{align}}\leftarrow\alpha\cdot\mathbb{1}[\hat{y}\neq y^{*}]\cdot\mathcal{C}+\beta\cdot\text{CE}(\hat{y},y^{*})$
+[…]
 7:if $\text{WMD}(\hat{y},y^{*})<\delta$ then
-8:     $h_{\text{pos}}\leftarrow\text{Embed}(\hat{y})$
+[…]
 9:end if
 10:if $\hat{y}\neq y^{*}\land\mathcal{C}>\tau_{1}\land\mathcal{U}<\tau_{2}$ then
-11:     $h_{\text{neg}}\leftarrow\text{Embed}(\hat{y})$
+[…]
 12:end if
 13:Compute $\mathcal{L}_{\text{contrast}}=\max(0,m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})+\text{sim}(h_{\text{anchor}},h_{\text{neg}}))$
-14:$\mathcal{L}_{\text{total}}\leftarrow\lambda_{1}\cdot\mathcal{L}_{\text{align}}+\lambda_{2}\cdot\mathcal{L}_{\text{contrast}}$
+[…]
 15:Update projection head using $\mathcal{L}_{\text{total}}$
 3 Methodology
-As discussed earlier, HonestVQA is a model-agnostic calibration framework designed to enhance ethical transparency in DocVQA systems.
+As discussed earlier, HonestVQA is a model-agnostic calibration framework designed to enhance ethical transparency in DocVQA models.
 It operates as a wrapper around pretrained DocVQA models (in our work, we evaluate our framework on top of pretrained models such as LayoutLMv3 Fujitake (2024), UDOP Wang et al.
@@
 Given a query-answer embedding $h_{\text{anchor}}$, we identify a positive sample $h_{\text{pos}}$ (semantically similar and ethically aligned) and a negative sample $h_{\text{neg}}$ (incorrect, overconfident, or potentially misleading).
-The contrastive loss is defined as according to Equation (3).
-$\mathcal{L}_{\text{contrast}}=$ $\max\Big{(}0,\ m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})$
+The contrastive loss is defined as according to Equation (4), where $\text{sim}(\cdot)$ denotes cosine similarity, and $m$ is a margin hyperparameter.
+$\mathcal{L}_{\text{contrast}}=\$$\max\Big(0,\ m-\text{sim}(h_{\text{anchor}},h_{\text{pos}})$
 $+\ \text{sim}(h_{\text{anchor}},h_{\text{neg}})\Big)$ (4)
-where $\text{sim}(\cdot)$ denotes cosine similarity, and $m$ is a margin hyperparameter.
 We use a projection head atop the DocVQA encoder to map answer embeddings into a low-dimensional calibrated honesty space.
@@
 $\text{WMD}(\hat{y},y^{*})<\delta$
-$\hat{y}\in\mathcal{A}_{\text{aligned}}\Longrightarrow$ semantically valid
+$\hat{y}\in\mathcal{A}_{\text{aligned}}\Longrightarrow\$semantically valid
 and agrees with ground truth. (6)
-Whereas, negative samples are drawn from high-confidence as shown in Equation (6).
-$\hat{y}_{\text{neg}}:$ $\mathbb{1}[\hat{y}_{\text{neg}}\neq y^{*}]$
+Whereas, negative samples are drawn from high-confidence as shown in Equation (7), where $\tau_{1}$ and $\tau_{2}$ are confidence and entropy thresholds, respectively.
+$\hat{y}_{\text{neg}}:\$$\mathbb{1}[\hat{y}_{\text{neg}}\neq y^{*}]$
 $\land\ \mathcal{C}(D,Q)>\tau_{1}$
 $\land\ \mathcal{U}(D,Q)<\tau_{2}$ (7)
-where $\tau_{1}$ and $\tau_{2}$ are confidence and entropy thresholds, respectively.
 3.4 Training Module
@@
 HonestVQA is not a moral arbiter but a mechanism for promoting caution and transparency in DocVQA behavior.
+Furthermore, our framework is fully multimodal.
+HonestVQA operates as a wrapper on LayoutLMv3, UDOP, and DONUT—all of which jointly model text + layout + vision.
+The uncertainty and contrastive modules explicitly leverage multimodal embeddings (e.g., attention-grounded text regions).
+This is confirmed by our IoU multimodal consistency evaluation (see Table 6), where HonestVQA achieves significantly higher alignment between visual attention and textual grounding compared to baselines.
 4 Experimental Setup
@@
 macro) to quantify answer correctness.
-To capture the alignment between model confidence and actual correctness, we introduce the Honesty Score (H-Score), which penalizes overconfident incorrect predictions while rewarding calibrated confidence on correct answers.
-Additionally, the Ethical Confidence Index (ECI) evaluates the model’s ability to appropriately express uncertainty, especially under ambiguous or insufficient information.
-Note: In the tables, color coding is used to indicate the performance of the models: Blue represents the best performance, Green indicates the second-best performance, and Yellow denotes the third-best performance.
-Whereas, ↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
+To capture the alignment between model confidence and actual correctness, we introduce the H-Score, which penalizes overconfident incorrect predictions while rewarding calibrated confidence on correct answers.
+Additionally, the ECI evaluates the model’s ability to appropriately express uncertainty, especially under ambiguous or insufficient information.
+Note: In the tables, Green indicate the best-performing scores.
+↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
 4.2.1 Theoretical Guarantees for Evaluation Metrics
-In this section, we provide formal lemmas and proofs to establish the theoretical soundness of the proposed Honesty Score (H-Score) and Ethical Confidence Index (ECI), which measure the alignment between model confidence, accuracy, and ethical transparency in DocVQA.
+In this section, we provide formal lemmas and proofs to establish the theoretical soundness of the proposed H-Score and ECI, which measure the alignment between model confidence, accuracy, and ethical transparency in DocVQA.
 Lemma 4.1 (Calibration Bound of Honesty Score).
-Let $C(D,Q)$ be the confidence score output by a DocVQA model for a given document-question pair $(D,Q)$, and let $A(D,Q)\in\{0,1\}$ be the corresponding accuracy indicator, where $1$ denotes a correct answer and 00 an incorrect one.
+Let $C(D,Q)$ be the confidence score output by a DocVQA model for a given document-question pair $(D,Q)$, and let $A(D,Q)\in\{0,1\}$ be the corresponding accuracy indicator, where $1$ denotes a correct answer and 0 an incorrect one.
 Assume that $C(D,Q)$ is bounded in $[0,1]$.
@@
 By definition, calibration error measures the absolute difference between predicted confidence and true correctness.
-Since $A(D,Q)$ is binary, the expectation of $|C-A|$ captures the average misalignment.
+Since $A(D,Q)$ is binary, the expectation of |CA||C-A| captures the average misalignment.
 Rearranging, $\mathrm{H}=1-\mathbb{E}[|C-A|]$.
@@
 DONUT Li et al. (2024) + HonestVQA 74.0 70.5 68.0 64.7 72.2 68.8
-Table 1: Answer correctness comparison on SpDocVQA, InfographicsVQA, and SROIE datasets
+Table 1: Performance comparison of base DocVQA models and their HonestVQA-enhanced versions across SpDocVQA, InfographicsVQA, and SROIE datasets, showing improvements in both Accuracy and Macro F1 scores.
 Model SpDocVQA InfographicsVQA SROIE
@@
 DONUT Li et al. (2024) + HonestVQA 0.120 0.139 0.125 0.143 0.122 0.141
-Table 2: Calibration metrics on SpDocVQA, InfographicsVQA, and SROIE datasets
+Table 2: Calibration performance of base DocVQA models and their HonestVQA-enhanced versions on SpDocVQA, InfographicsVQA, and SROIE datasets, measured by H-Score and ECI.
+Lower values indicate better model calibration and reduced overconfidence.
 4.3 Hyperparameters
@@
 For sample selection in the contrastive module, the WMD threshold $\delta$ is set to 0.4, while the confidence and entropy thresholds, $\tau_{1}$ and $\tau_{2}$, are chosen as 0.8 and 0.5, respectively, to effectively identify semantically valid positive samples and high-confidence misleading negatives.
-Note: We will make the code publicly available post-acceptance.
+Note: At inference, the model does not access ground-truth answers.
+Instead, it relies on the learned calibration from training–the confidence score $C(D,Q)$ and entropy $U(D,Q)$ are computed from the predicted distribution to identify uncertain or overconfident answers.
+The model can then suppress or abstain from outputs under high uncertainty, while the contrastive embeddings help separate ethically aligned responses from potentially misleading ones.
+Train Domain Test Domain Model Accuracy (%) ↑ Macro F1 (%) ↑
+SpDocVQA InfographicsVQA HonestVQA 74.2 71.8
+InfographicsVQA SpDocVQA HonestVQA 78.9 76.1
+SROIE InfographicsVQA HonestVQA 70.5 67.2
+SpDocVQA SROIE HonestVQA 72.6 69.8
+InfographicsVQA SROIE HonestVQA 73.1 70.4
+SROIE SpDocVQA HonestVQA 75.0 72.3
+Table 3: Cross-domain performance of HonestVQA, showing Accuracy and Macro F1 when trained on one dataset and evaluated on another.
+Results indicate robust generalization across SpDocVQA, InfographicsVQA, and SROIE datasets.
+Configuration SpDocVQA InfographicsVQA SROIE
+Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓
+Full HonestVQA Model 75.9 0.113 68.3 0.134 80.2 0.096
+No Alignment Loss 72.1 0.172 65.0 0.193 76.4 0.141
+No Contrastive Loss 73.0 0.160 65.2 0.189 77.2 0.133
+Table 4: Ablation study of HonestVQA showing the impact of removing alignment or contrastive loss on Accuracy and H-Score across SpDocVQA, InfographicsVQA, and SROIE.
+The full model consistently achieves the best performance and calibration.
 5 Experimental Analysis
@@
 Table 1 presents these results on the three datasets.
-It is evident that HonestVQA consistently improves accuracy by approximately 3 to 4 percentage points and macro F1-score by nearly 4 points across all base models.
+It is evident that HonestVQA consistently improves accuracy by approximately 3% to 4% and macro F1-score by nearly 4% across all base models.
 For instance, LayoutLMv3 Fujitake (2024) improves from 72.3% to 75.9% in accuracy and from 68.5% to 72.8% in macro F1-score.
@@
 This demonstrates that HonestVQA effectively mitigates the risk of misleading the user by suppressing confident but incorrect answers.
-Train Domain Test Domain Model Accuracy (%) ↑ Macro F1 (%) ↑
-SpDocVQA InfographicsVQA HonestVQA 74.2 71.8
-InfographicsVQA SpDocVQA HonestVQA 78.9 76.1
-SROIE InfographicsVQA HonestVQA 70.5 67.2
-SpDocVQA SROIE HonestVQA 72.6 69.8
-InfographicsVQA SROIE HonestVQA 73.1 70.4
-SROIE SpDocVQA HonestVQA 75.0 72.3
-Table 3: Cross-domain hallucination detection performance of HonestVQA
-Configuration SpDocVQA InfographicsVQA SROIE
-Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓ Accuracy (%) ↑ H-Score ↓
-Full HonestVQA Model 75.9 0.113 68.3 0.134 80.2 0.096
-No Alignment Loss 72.1 0.172 65.0 0.193 76.4 0.141
-No Contrastive Loss 73.0 0.160 65.2 0.189 77.2 0.133
-Table 4: Ablation results across three datasets showing the impact of disabling individual components of HonestVQA
 5.2 Cross-Domain Generalization Testing
@@
 Similarly, the SROIE $\rightarrow$ SpDocVQA setup results in an F1 of 72.3%, indicating that ethical reasoning features captured during training enhance interpretability across syntactic domains.
-However, the model exhibits relatively lower performance when transferring from SpDocVQA to SROIE (F1 = 69.8%), highlighting the challenges posed by ethical hallucination detection under unfamiliar structural constraints.
+However, the model exhibits relatively lower performance when transferring from SpDocVQA $\rightarrow$ SROIE (F1 = 69.8%), highlighting the challenges posed by ethical hallucination detection under unfamiliar structural constraints.
 Nonetheless, the use of uncertainty-aware calibration via confidence-alignment and contrastive ethical loss contributes to soft regularization of decision boundaries, allowing for improved generalization even in low-overlap semantic settings.
@@
 This supports our hypothesis that the inclusion of contrastive ethical supervision enforces more generalizable representations.
-Furthermore, the relative drop in performance in domain transfer settings (typically within 4–6% of in-domain results) underscores the importance of calibration-aware models in mitigating performance degradation due to domain shift.
+Furthermore, the relative drop in performance in domain transfer settings (typically within 4%–6% of in-domain results) underscores the importance of calibration-aware models in mitigating performance degradation due to domain shift.
 Hyperparameter Values Tested SpDocVQA (Acc ↑, H ↓) InfographicsVQA (Acc ↑, H ↓) SROIE (Acc ↑, H ↓)
@@
 0.10 (73.1, 0.142) (65.4, 0.172) (77.3, 0.125)
-Table 5: Sensitivity analysis of alignment and contrastive hyperparameters on SpDocVQA, InfographicsVQA, and SROIE.
-Note: Accuracy (Acc) and H-Score (H) are used as abbreviated.
-Bold rows indicate optimal settings.
+Table 5: Sensitivity analysis of alignment and contrastive hyperparameters on SpDocVQA, InfographicsVQA, and SROIE datasets. Note: Accuracy (Acc) and H-Score (H) are used as abbreviated.
 Model SpDocVQA InfographicsVQA SROIE
@@
 6.2 Computational Analysis
-Practical deployment of DocVQA systems, particularly those incorporating hallucination detection like HonestVQA, requires a thorough understanding of computational costs and resource demands.
-To assess the efficiency of HonestVQA relative to baseline models, we conduct a detailed latency and resource usage analysis focusing on inference time, Floating-Point Operations (FLOPs), and memory consumption.
-All models are tested on identical hardware setups: an NVIDIA RTX 3090 GPU and an Intel Xeon CPU, with batch size fixed to 1 to simulate single-query inference typical of interactive systems.
-Table 7 summarizes the key efficiency metrics averaged over 1000 inference runs per dataset.
-From Table 7, HonestVQA demonstrates an average inference latency of approximately 119.6 ms per query, which is about 6%-20% slower than the baseline models due to the additional uncertainty calibration and contrastive enforcement modules integrated into the architecture.
-Despite this, the latency remains within acceptable limits for real-time applications, especially in contexts where reliability and ethical correctness outweigh marginal speed gains.
-Regarding computational complexity, HonestVQA requires on average 69.4 GFLOPs, representing a 7%-18% increase relative to baselines.
-This overhead is primarily attributable to the dual loss functions enforcing confidence-accuracy alignment and ethical contrastive constraints, which necessitate additional forward and backward computations during inference calibration.
-Memory usage also increases modestly to 3075 MB on average, about 5%-14% more than competing models.
-The memory footprint remains manageable on modern GPUs, supporting the feasibility of deploying HonestVQA on standard deep learning infrastructure without requiring specialized hardware.
+We evaluate the efficiency of HonestVQA against baseline DocVQA models in terms of inference latency, FLOPs, and memory usage.
+All models are tested using an NVIDIA RTX 3090 GPU and Intel Xeon CPU with a batch size of 1 to simulate real-time settings.
+As shown in Table 7, HonestVQA incurs an average latency of 119.6 ms per query—6%–20% slower than baselines—due to uncertainty calibration and contrastive modules.
+It consumes 69.4 GFLOPs (7%–18% higher) and 3075 MB memory (5%–14% higher).
+Despite the overhead, it remains deployable in real-time systems where ethical reliability is critical.
+(a) Ethical Risk Heatmap Confidence vs Uncertainty
+(b) Semantic Drift under Ambiguity
+(c) Contrastive Embedding Separation
+(d) Attention IoU over Epochs
+Figure 1: HonestVQA enhances ethical calibration, semantic stability, embedding separation, and multimodal grounding through uncertainty-aware learning,
+6.3 Qualitative Analysis
+HonestVQA improves ethical alignment by reducing overconfidence in uncertain scenarios, as seen in the risk heatmap (Fig.
+1a).
+Semantic drift under ambiguity is mitigated, with more stable embeddings (Fig.
+1b).
+Contrastive embedding separation (Fig.
+1c) shows clearer distinction between aligned and misaligned responses, supporting improved representation learning.
+Finally, Fig.
+1d shows consistent attention patterns over epochs, highlighting better multimodal grounding and interpretability.
 7 Conclusion and Future Works
-In this work, we introduced HonestVQA, a novel framework that integrates uncertainty-aware alignment and contrastive ethical enforcement to effectively detect hallucinations in DocVQA systems.
+In this work, we introduced HonestVQA, a novel framework that integrates uncertainty-aware alignment and contrastive ethical enforcement to effectively detect hallucinations in DocVQA models.
 Through comprehensive experiments on diverse datasets—we demonstrated significant improvements in answer correctness, calibration, and cross-domain generalization compared to strong baselines.
 Our ablation studies confirmed the complementary role of each component, and efficiency analyses showed HonestVQA practical feasibility.
-Future research will focus on advancing domain adaptation techniques to further enhance robustness across unseen data distributions, and exploring lightweight calibration modules for deployment on edge devices.
+Future research will focus on advancing domain adaptation models to further enhance robustness across unseen data distributions, and exploring lightweight calibration modules for deployment on edge devices.
 Additionally, incorporating user feedback for interactive hallucination correction and extending the framework to multimodal dialogue systems represent promising directions to improve DocVQA reliability and ethical safety.
@@
 Ethics Statement
-This work aims to enhance the trustworthiness and ethical reliability of DocVQA systems by reducing hallucinated and potentially misleading answers.
+This work aims to enhance the trustworthiness and ethical reliability of DocVQA models by reducing hallucinated and potentially misleading answers.
 HonestVQA promotes transparency through uncertainty-aware calibration, encouraging responsible AI deployment.
@@
 All datasets used comply with their respective licenses, and no private or sensitive data was involved.
-We encourage further research on fairness, bias mitigation, and inclusivity to ensure equitable AI systems, and advocate for ongoing monitoring of model outputs to safeguard against misuse or unintended harm.
+We encourage further research on fairness, bias mitigation, and inclusivity to ensure equitable AI models, and advocate for ongoing monitoring of model outputs to safeguard against misuse or unintended harm.
+Acknowledgments
+This research was supported by Jamia Hamdard University, New Delhi, India and funded through the Visvesvaraya Fellowship, Government of India.
+Furthermore, this research is partially sponsored by funding from Yonghua Foundation.